The hypotenuse of a right triangle is 13cm. The sum of the lengths of the two other sides is 17cm. Find the unknown lengths.
Answer by ankor@dixie-net.com(22740)

The hypotenuse of a right triangle is 13cm.
The sum of the lengths of the two other sides is 17cm.
Find the unknown lengths.
:
Let x = length of one side
then
(17-x) = length of the other side
;
Using pythag
x^2 + (17-x)^2 = 13^2
FOIL
x^2 + (289 - 34x + x^2) = 169
Which is
2x^2 - 34x + 289 - 169 = 0
:
2x^2 - 34x + 120 = 0
Simplify, divide by 2
x^2 - 17x + 60 = 0
Factor
(x-12)(x-5) = 0
x = 12, the other side is 5
or
x = 5, the other side is 12